29.50(1096-n) + 35.40n = 37
Found 2 solutions by mukhopadhyay, checkley71:
Answer by mukhopadhyay(490)

29.50(1096-n) + 35.40n = 37
=>32332 + n(35.4-29.5) = 37
=>5.9n = -32995
=>n = -5592.3729 (approx)

Answer by checkley71(8403)

29.5(1096-N)+35.4N=37 OR 32332-29.5N+35.4N=37 OR 35.4N-29.5N=-32332+37 OR
5.9N=-32295 OR N=-32295/5.9 OR N=-5473.7288
PROOF 32332-29.5*-5473.7288+35.4*-5473.7288=37 OR
32332+161475-193770=37 OR 193807-193770=37 OR 37=37